OpenStudy (anonymous):

A ladder leans against a house at a 60° angle to the ground. If the ladder extends to a length of 166 inches, what is the height of the house to the nearest hundredth of an inch? House with a ladder measuring 166 inches forming a right triangle. The base angle formed by the ladder is 60 degrees. A ladder leans against a house at a 60° angle to the ground. If the ladder extends to a length of 166 inches, what is the height of the house to the nearest hundredth of an inch? House with a ladder measuring 166 inches forming a right triangle. The base angle formed by the ladder is 60 degrees. @Mathematics

OpenStudy (anonymous):



OpenStudy (anonymous):

Sin(60) = h/166 So, height of house = 166 * sin(60)